# Quantum Channel Testing in Average-Case Distance

### v1

@@ 1.3 Channel certification and tomography in ACID distance
The case of Theorem 1.8 where $\mathcal{N}$ is the identity channel on $\mathbb{C}^{d\times d}$ is particularly interesting. By the Fuchs–van de Graaf inequalities, the ACID distance $\frac{1}{2}{\|\mathcal{M}-\mathcal{N}\|}_{J}=\frac{1}{2}{\|J_{\mathcal{M}}-\Phi_{d}\|}_{1}$ is closely related to the entanglement fidelity [Wil19, Definition 9.5.1] $\operatorname*{tr}\mathopen{}\mathclose{{}\left(J_{\mathcal{M}}\Phi_{d}}\right)$ between $\mathcal{M}$ and the identity channel with respect to the maximally entangled state. Fawzi et al. [FFGO23, Lemma A.1] proved that if $d$ is large, then $\operatorname*{tr}\mathopen{}\mathclose{{}\left(J_{\mathcal{M}}\Phi_{d}}\right)$ is a close approximation of $\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[\operatorname*{tr}(\mathcal{M}(\boldsymbol{\psi})\boldsymbol{\psi})}\right]$ (where $\boldsymbol{\psi}$ is Haar random), a quantity which is a standard measure for quantifying errors in physical implementations of quantum gates [KLDF16, Eq. 1].
@@
Theorem 1.8 does not generalize to arbitrary channels $\mathcal{N}$ however. For example, let $\mathcal{N}$ be the channel that replaces its input with a known state $\sigma$ (i.e. $\mathcal{N}(X)=\operatorname*{tr}(X)\sigma$), and suppose that the blackbox channel $\mathcal{M}$ is promised to replace its input with some unknown state $\rho$ (i.e. $\mathcal{M}(X)=\operatorname*{tr}(X)\rho$). It is straightforward to verify that ${\|\mathcal{M}-\mathcal{N}\|}_{J}={\|\rho-\sigma\|}_{1}$, and query access to $\mathcal{M}$ is equivalent to sample access to $\rho$, so testing identity to $\mathcal{N}$ in ACID distance is no easier than testing identity to $\sigma$ in trace distance. If $\sigma$ is the maximally mixed state for example, i.e. if $\mathcal{N}$ is the completely depolarizing channel, then the complexity of this task is $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{out}}^{3/2}/\varepsilon^{2}}\right)$ in the ancilla-free, adaptive model [CLHL22, Theorem 6.1]. This is why we specifically considered pure state replacement channels in Theorem 1.8.
@@
The $\Omega(d_{\mathrm{in}}/\varepsilon^{2})$ lower bound from Theorem 1.9 matches the dependence on $d_{\mathrm{in}}$ and $\varepsilon$ from Theorem 1.6 in the same query model, and along with the above discussion implies an $\Omega\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}+d_{\mathrm{out}}^{3/2}}\right)\big{/}\varepsilon^{2}}\right)$ lower bound for testing identity to the completely depolarizing channel in this model. We conjecture that this lower bound can be improved to $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}d_{\mathrm{out}}^{3/2}\big{/}\varepsilon^{2}}\right)$.
We also briefly consider a nonstandard query model, where it turns out that channel certification can always be done with complexity independent of the input dimension. King, Wan and McClean [KWM24] proposed a model of quantum state testing where copies of both $\rho$ and $\rho^{\top}$ are provided, and gave several examples [KWM24, Appendix D] where this may be a physically realistic assumption. Analogously, for a channel $\mathcal{M}$ we define ¯$\overline{\mathcal{M}}(X)\coloneqq M\mathopen{}\mathclose{{}\left(X^{\top}}\right)^{\top}$. The fact that $\overline{\mathcal{M}}$ is a channel is most easily seen by considering its Kraus decomposition (see Eq. 4), which also makes it clear that $\overline{\mathcal{M}}$ is the element-wise complex conjugate of $\mathcal{M}$. For example, if a channel $\mathcal{M}$ is defined by evolving a real-valued Hamiltonian forward in time, then $\overline{\mathcal{M}}$ is defined by evolving that same Hamiltonian backward in time. If $\mathcal{M}$ is implemented by a quantum circuit over the gate set $\{H,T,\mathrm{Toffoli}\}$, then $\overline{\mathcal{M}}$ can be implemented by substituting $T^{\dagger}$ for $T$ throughout the circuit. [^10] However, if our motivation is to test whether an alleged circuit implementation of $\mathcal{N}$ is accurate, then there is no guarantee that faulty implementations of $\mathcal{N}$ and $\overline{\mathcal{N}}$ would be $\mathcal{M}$ and $\overline{\mathcal{M}}$ respectively for the same channel $\mathcal{M}$. We prove the following:
@@
Finally we consider the complexity of channel tomography in ACID distance, as a benchmark against which to compare our results about channel testing (as testing is trivially no more difficult than tomography). O’Donnell and Wright [OW15, Theorem 1.10] gave an algorithm that performs an entangled measurement on $O(d^{2}/\varepsilon^{2})$ copies of a density matrix $\rho\in\mathbb{C}^{d\times d}$, and with probability at least $2/3$ outputs the description of a density matrix $\sigma\in\mathbb{C}^{d\times d}$ such that ${\|\rho-\sigma\|}_{1}\leq\varepsilon$. Specializing to the case $\rho=J_{\mathcal{M}}$ immediately implies the following:
@@
We prove a nontrivial matching lower bound for fixed $\varepsilon$, even for adaptive algorithms:
@@
Since the ACID norm is trivially at most the diamond norm, the following results of Oufkir [O23] resolve the complexity of incoherent, non-adaptive channel tomography in both ACID and diamond distances:
Theorem 1.13 (Incoherent channel tomography [O23, Theorems 3.3 and 2.1111111The lower bound is stated in terms of the diamond norm, but inspection of the proof reveals that it goes through the ACID norm.]).
There is an ancilla-free, non-adaptive algorithm that makes $\tilde{O}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{3}d_{\mathrm{out}}^{3}\big{/}\varepsilon^{2}}\right)$ queries to a channel $\mathcal{M}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$, and outputs the description of a channel $\mathcal{N}$ such that ${\|\mathcal{M}-\mathcal{N}\|}_{\diamond}\leq\varepsilon$ with probability at least $2/3$. Furthermore $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{3}d_{\mathrm{out}}^{3}\big{/}\varepsilon^{2}}\right)$ queries are necessary for this task when $d_{\mathrm{out}}\geq 4$, even using ancillae (but not coherence or adaptivity) and even with the ACID norm in place of the diamond norm.
@@
Unitary channel $\diamond$ $O(d/\varepsilon^{4})$ [FFGO23, Thm. 3.1] *$\Omega\mathopen{}\mathclose{{}\left(d/\varepsilon^{2}}\right)$ [FFGO23, Thm. 3.1] *$\Omega\mathopen{}\mathclose{{}\left(d^{1/2}}\right)$ [MdW13, Sec. 5.1.1]
@@
$J$ $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{3}d_{\mathrm{out}}^{3}\big{/}\varepsilon^{2}}\right)$ Theorem 1.13 $O\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{2}/\varepsilon^{2}}\right)$ Theorem 1.11 *$\boldsymbol{\tilde{\Omega}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{2}}\right)}$ Theorem 1.12
@@ 1.4 Open problems
The sample complexity of testing identity to a fixed state $\sigma\in\mathbb{C}^{d\times d}$ using unentangled measurements is roughly $d^{3/2}/\varepsilon^{2}$ times the (square) fidelity of $\sigma$ with the maximally mixed state [CLO22, CLHL22]. Analogously, what is the query complexity of testing identity to a fixed channel $\mathcal{N}$ in any of the query models that we have discussed? One may approach this question by trying to close some of the gaps between the upper and lower bounds in Table 1. What if we consider tolerant testing, where the goal is to decide whether ${\|\mathcal{M}-\mathcal{N}\|}_{J}\leq\delta$ or ${\|\mathcal{M}-\mathcal{N}\|}_{J}\geq\varepsilon$?
Unitary tomography
Haah, Kothari, O’Donnell and Tang [HKOT23] proved that $\Theta(d^{2}/\varepsilon)$ queries are necessary and sufficient to learn a $d$-dimensional unitary channel in diamond distance [^12] Specifically, their upper bound is ancilla-free and adaptive, and their lower bound is against ancilla-assisted, coherent, adaptive algorithms., which is a polynomial-factor improvement on the complexity of generic channel tomography in diamond distance (cf. Theorem 1.13 for example). Does a similar speedup hold for unitary tomography in ACID distance?
@@
A $k$-junta is a channel from $\mathopen{}\mathclose{{}\left(\mathbb{C}^{2\times 2}}\right)^{\otimes n}$ to $\mathopen{}\mathclose{{}\left(\mathbb{C}^{2\times 2}}\right)^{\otimes n}$ that acts nontrivially on at most $k$ qubits. Chen, Nadimpalli and Yuen [Chen2023Junta] proved that $\tilde{\Theta}(\sqrt{k})$ queries to a unitary channel are necessary and sufficient to test whether it is a $k$-junta or far from all $k$-juntas, and that $\tilde{\Theta}(4^{k})$ queries are necessary and sufficient to learn a unitary $k$-junta. (We have suppressed the dependence on $\varepsilon$ for simplicity.) In Section 4.2 we show that their distance measure is proportional to ACID distance, so it is natural to ask whether their results generalize to the case where the blackbox channel is not necessarily unitary, with distance measured in ACID distance. Bao and Yao [Bao23Junta] proved similar results (except with only an $\tilde{O}(k)$ upper bound for testing) when the blackbox channel is not necessarily unitary, but they measured distance between channels by the $\ell_{2}$ distance between their Choi states, a quantity which is only loosely related to the ACID norm via Cauchy-Schwarz and the fact that the 2-norm is at most the 1-norm.
@@ 2.1 Quantum states and transformations
We denote a superoperator $\mathcal{L}$ applied to an input $X$ by any of $\mathcal{L}(X)$ or $\mathcal{L}\cdot X$ or $\mathcal{L}X$ [^13] No relation to the Pauli $X$ matrix. If we define a superoperator $\mathcal{L}\in\mathsf{S}(d,*)$ by its action on a matrix $X$ without any further context, then $X$ implicitly ranges over all matrices in $\mathbb{C}^{d\times d}$. We write superoperators in mathcal font.
@@
A register is a finite-dimensional complex Hilbert space. We write $\mathsf{AB}$ to denote the tensor product of registers $\mathsf{A}$ and $\mathsf{B}$, and $\mathsf{D}(\mathsf{A})$ to denote the set of density matrices in a register $\mathsf{A}$. We also write $\mathsf{D}(d_{1}\otimes d_{2})$ to denote the set of density matrices in $\mathbb{C}^{d_{1}\times d_{1}}\otimes\mathbb{C}^{d_{2}\times d_{2}}$, and similarly for the other notation above.
@@ 2.2 Matrix norms and fidelity
The (square) fidelity of two density matrices $\rho$ and $\sigma$ is the quantity $\mathrm{F}\mathopen{}\mathclose{{}\left(\rho,\sigma}\right)\coloneqq{\|\sqrt{\rho}\sqrt{\sigma}\|}_{1}^{2}$. In particular, if $\rho\in\mathsf{D}(d)$ is an arbitrary density matrix then $\mathrm{F}\mathopen{}\mathclose{{}\left(\rho,I/d}\right)=\operatorname*{tr}(\sqrt{\rho})^{2}/d$, and if furthermore $\psi\in\mathsf{D}(d)$ is a pure state then $\mathrm{F}\mathopen{}\mathclose{{}\left(\rho,\psi}\right)=\operatorname*{tr}(\rho\psi)$. We also use the following half of the Fuchs–-van de Graaf inequalities: for all density matrices $\rho$ and $\sigma$,

### v2

@@ 1.3 Channel certification and tomography in ACID distance
The case of Theorem 1.8 where $\mathcal{N}$ is the identity channel on $\mathbb{C}^{d\times d}$ is particularly interesting. By the Fuchs–van de Graaf inequalities, the ACID distance $\frac{1}{2}{\|\mathcal{M}-\mathcal{N}\|}_{J}=\frac{1}{2}{\|J_{\mathcal{M}}-\Phi_{d}\|}_{1}$ is quadratically related to the entanglement fidelity [Wil19, Definition 9.5.1] $\operatorname*{tr}\mathopen{}\mathclose{{}\left(J_{\mathcal{M}}\Phi_{d}}\right)$ between $\mathcal{M}$ and the identity channel with respect to the maximally entangled state. Fawzi et al. [FFGO23, Lemma A.1] proved that if $d$ is large, then $\operatorname*{tr}\mathopen{}\mathclose{{}\left(J_{\mathcal{M}}\Phi_{d}}\right)$ is a close approximation of $\operatorname*{\mathbb{E}}\mathopen{}\mathclose{{}\left[\operatorname*{tr}(\mathcal{M}(\boldsymbol{\psi})\boldsymbol{\psi})}\right]$ (where $\boldsymbol{\psi}$ is Haar random), a quantity which is a standard measure for quantifying errors in physical implementations of quantum gates [KLDF16, Eq. 1].
@@
Theorem 1.8 does not generalize to arbitrary channels $\mathcal{N}$ however. For example, let $\mathcal{N}$ be the channel that replaces its input with a known state $\sigma$ (i.e. $\mathcal{N}(X)=\operatorname*{tr}(X)\sigma$), and suppose that $\mathcal{M}$ is promised to replace its input with some unknown state $\rho$ (i.e. $\mathcal{M}(X)=\operatorname*{tr}(X)\rho$). It is straightforward to verify that ${\|\mathcal{M}-\mathcal{N}\|}_{J}={\|\rho-\sigma\|}_{1}$, and query access to $\mathcal{M}$ is equivalent to sample access to $\rho$, so testing identity to $\mathcal{N}$ in ACID distance is no easier than testing identity to $\sigma$ in trace distance. If $\sigma$ is the maximally mixed state for example, i.e. if $\mathcal{N}$ is the completely depolarizing channel, then this requires $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{out}}^{3/2}/\varepsilon^{2}}\right)$ queries in the ancilla-free, adaptive model [CLHL22, Theorem 6.1]. This is why we specifically considered pure state replacement channels in Theorem 1.8.
@@
The $\Omega(d_{\mathrm{in}}/\varepsilon^{2})$ lower bound from Theorem 1.9 matches the dependence on $d_{\mathrm{in}}$ and $\varepsilon$ from Theorem 1.6 in the same query model, and along with the above discussion implies an $\Omega\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}+d_{\mathrm{out}}^{3/2}}\right)\big{/}\varepsilon^{2}}\right)$ lower bound for testing identity to the completely depolarizing channel in this model. We conjecture that this lower bound can be improved to $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}d_{\mathrm{out}}^{3/2}\big{/}\varepsilon^{2}}\right)$, which would match our upper bound.
We also briefly consider a nonstandard query model, where it turns out that channel certification can always be done with complexity independent of the input dimension. King, Wan and McClean [KWM24] proposed a model of quantum state testing with sample access to both $\rho$ and $\rho^{\top}$, and gave several examples [KWM24, Appendix D] where this may be a physically realistic assumption. Analogously, for a channel $\mathcal{M}$ we define ¯$\overline{\mathcal{M}}(X)\coloneqq M\mathopen{}\mathclose{{}\left(X^{\top}}\right)^{\top}$. The fact that $\overline{\mathcal{M}}$ is a channel is most easily seen by considering its Kraus decomposition (see Eq. 3), which also illustrates that $\overline{\mathcal{M}}$ is the element-wise complex conjugate of $\mathcal{M}$. For example, if $\mathcal{M}$ is defined by evolving a real-valued Hamiltonian forward in time, then $\overline{\mathcal{M}}$ is defined by evolving that same Hamiltonian backward in time. If $\mathcal{M}$ is implemented by a quantum circuit over the gate set $\{H,T,\mathrm{Toffoli}\}$, then $\overline{\mathcal{M}}$ can be implemented by substituting $T^{\dagger}$ for $T$ throughout that circuit. [^10] However, if our motivation is to test whether an alleged circuit implementation of $\mathcal{N}$ is accurate, then there is no guarantee that faulty implementations of $\mathcal{N}$ and $\overline{\mathcal{N}}$ would be $\mathcal{M}$ and $\overline{\mathcal{M}}$ respectively for the same channel $\mathcal{M}$. We prove the following:
@@
Finally we consider the complexity of channel tomography in ACID distance, as a benchmark against which to compare our results about channel testing (as testing trivially reduces to tomography). We prove the following by a blackbox reduction to state tomography on $J_{\mathcal{M}}$, followed by post-processing to ensure that the output is a channel:
@@
We also prove a nontrivial matching lower bound for fixed $\varepsilon$, even for adaptive algorithms:
@@
Since the ACID norm is trivially at most the diamond norm, results of Oufkir [O23] resolve the complexity of incoherent, non-adaptive channel tomography in both ACID and diamond distances:
Theorem 1.13 (Incoherent channel tomography [O23, Theorems 3.3 and 2.1111111The lower bound is stated in terms of diamond distance, but inspection of the proof reveals that it holds for ACID distance.]).
There is an ancilla-free, non-adaptive algorithm that makes $\tilde{O}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{3}d_{\mathrm{out}}^{3}\big{/}\varepsilon^{2}}\right)$ queries to a channel $\mathcal{M}:\mathbb{C}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}\to\mathbb{C}^{d_{\mathrm{out}}\times d_{\mathrm{out}}}$, and outputs the description of a channel $\mathcal{N}$ such that ${\|\mathcal{M}-\mathcal{N}\|}_{\diamond}\leq\varepsilon$ with probability at least $2/3$. Furthermore $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{3}d_{\mathrm{out}}^{3}\big{/}\varepsilon^{2}}\right)$ queries are necessary for this task when $d_{\mathrm{out}}\geq 4$, even using ancillae (but not coherence or adaptivity) and with the ACID norm in place of the diamond norm.
Similarly, the following upper bound of Haah, Kothari, O’Donnell and Tang [HKOT23, Theorem 1.1] and lower bound of Zhao, Lewis, Kannan, Quek, Huang and Caro [Zhao+23, $G=d^{2}$ case of Theorem 4] [^12] In Section 4.2 we explain why Zhao et al.’s distance is equivalent to ACID distance. resolve the complexity of unitary tomography in both ACID and diamond distances:
Theorem 1.14 (Unitary tomography [HKOT23, Zhao+23]).
There is an ancilla-free, adaptive algorithm that makes $O(d^{2}/\varepsilon)$ queries to a unitary channel $\mathcal{M}:\mathbb{C}^{d\times d}\to\mathbb{C}^{d\times d}$, and outputs the description of a unitary channel $\mathcal{N}$ such that ${\|\mathcal{M}-\mathcal{N}\|}_{\diamond}\leq\varepsilon$ with probability at least $2/3$. Furthermore $\Omega(d^{2}/\varepsilon)$ queries are necessary for this task, even using ancillae and coherence, and even with the ACID norm in place of the diamond norm.
@@
Unitary channel $\diamond$ $O(d/\varepsilon^{4})$ [FFGO23, Thm. 3.1] *$\Omega\mathopen{}\mathclose{{}\left(d/\varepsilon^{2}}\right)$ [FFGO23, Thm. 3.1]
@@
$J$ $\Omega\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{3}d_{\mathrm{out}}^{3}\big{/}\varepsilon^{2}}\right)$ Theorem 1.13 $\boldsymbol{O\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{2}/\varepsilon^{2}}\right)}$ Theorem 1.11 *$\boldsymbol{\tilde{\Omega}\mathopen{}\mathclose{{}\left(d_{\mathrm{in}}^{2}d_{\mathrm{out}}^{2}}\right)}$ Theorem 1.12
@@ 1.4 Open problems
The sample complexity of testing identity to a fixed state $\sigma\in\mathbb{C}^{d\times d}$ using unentangled measurements is roughly $d^{3/2}/\varepsilon^{2}$ times the (square) fidelity of $\sigma$ with the maximally mixed state [CLO22, CLHL22]. Analogously, what is the query complexity of testing identity to a fixed channel $\mathcal{N}$ in any of the query models that we have discussed? One may approach this question by trying to close some of the gaps between the upper and lower bounds in Table 1. What if we consider tolerant testing, where the goal is to decide whether ${\|\mathcal{M}-\mathcal{N}\|}_{J}\leq\delta$ or ${\|\mathcal{M}-\mathcal{N}\|}_{J}\geq\varepsilon$? What if we also require our protocols to be computationally efficient, for example by sampling states from a locally scrambled ensemble [Zhao+23, Definition 1] instead of the Haar measure?
Testing and tomography of channels with bounded gate complexity
Zhao et al. [Zhao+23, Theorem 4] proved that $\tilde{O}\mathopen{}\mathclose{{}\left(G/\varepsilon\cdot\min\mathopen{}\mathclose{{}\left(1/\varepsilon,\sqrt{d}}\right)}\right)$ queries suffice and $\Omega(G/\varepsilon)$ queries are necessary to learn in ACID distance a $d$-dimensional unitary channel comprised of $G$ two-qubit gates. Does a similar statement hold for arbitrary channels? What about for testing rather than tomography?
@@
A $k$-junta is a channel from $\mathopen{}\mathclose{{}\left(\mathbb{C}^{2\times 2}}\right)^{\otimes n}$ to $\mathopen{}\mathclose{{}\left(\mathbb{C}^{2\times 2}}\right)^{\otimes n}$ that acts nontrivially on at most $k$ qubits. Chen, Nadimpalli and Yuen [Chen2023Junta] proved that $\tilde{\Theta}(\sqrt{k})$ queries to a unitary channel are necessary and sufficient to test whether it is a $k$-junta or far from all $k$-juntas, and that $\tilde{\Theta}(4^{k})$ queries are necessary and sufficient to learn a unitary $k$-junta. (We have suppressed the dependence on $\varepsilon$ for simplicity.) In Section 4.2 we show that their distance is proportional to ACID distance, so it is natural to ask whether their results generalize to the case where the blackbox channel is not necessarily unitary, with distance measured in the ACID norm. Bao and Yao [Bao23Junta] proved similar results (except with only an $\tilde{O}(k)$ upper bound for testing) when the blackbox channel is not necessarily unitary, but they measured distance between channels by the $\ell_{2}$ distance between their Choi states, a quantity which is only loosely related to the ACID norm via Cauchy-Schwarz and the fact that the 2-norm is at most the 1-norm.
@@ 2.1 Quantum states and transformations
We denote a superoperator $\mathcal{L}$ applied to an input $X$ by any of $\mathcal{L}(X)$ or $\mathcal{L}\cdot X$ or $\mathcal{L}X$ [^13] No relation to the Pauli $X$ matrix. If we define a superoperator $\mathcal{L}\in\mathsf{S}(d,*)$ by its action on an unspecified matrix $X$, then $X$ implicitly ranges over all matrices in $\mathbb{C}^{d\times d}$. We write superoperators in mathcal font.
@@
A register is a finite-dimensional complex Hilbert space. We write $\mathsf{AB}$ to denote the tensor product of registers $\mathsf{A}$ and $\mathsf{B}$, and $\mathsf{D}(\mathsf{A})$ to denote the set of density matrices in a register $\mathsf{A}$. We also write $\mathsf{D}(d_{1}\otimes d_{2})$ to denote the set of density matrices in $\mathbb{C}^{d_{1}\times d_{1}}\otimes\mathbb{C}^{d_{2}\times d_{2}}$, and similarly for $\mathsf{S}(\cdot,\cdot)$ and $\mathsf{C}(\cdot,\cdot)$.
@@ 2.2 Matrix norms and fidelity
The (square) fidelity of density matrices $\rho$ and $\sigma$ is the quantity $\mathrm{F}\mathopen{}\mathclose{{}\left(\rho,\sigma}\right)\coloneqq{\|\sqrt{\rho}\sqrt{\sigma}\|}_{1}^{2}$. In particular, if $\rho\in\mathsf{D}(d)$ is an arbitrary density matrix then $\mathrm{F}\mathopen{}\mathclose{{}\left(\rho,I/d}\right)=\operatorname*{tr}(\sqrt{\rho})^{2}/d$, and if furthermore $\psi\in\mathsf{D}(d)$ is a pure state then $\mathrm{F}\mathopen{}\mathclose{{}\left(\rho,\psi}\right)=\operatorname*{tr}(\rho\psi)$. We also use the following half of the Fuchs–-van de Graaf inequalities: for all density matrices $\rho$ and $\sigma$,